An artillery is pointed upward at an angle of 30 degrees with respect to the horizontal with
a muzzle velocity of 1500 ft/s. Ignoring the effect of air resistance, to what height will the
projectile rise and what will be its range?
1
Expert's answer
2020-10-12T07:50:49-0400

We can find the maximum height that the projectile can reach from the kinematic equation:


"v_{y,f}^2=v_{y,i}^2 - 2ay_{max},"

here, "v_{y,f} = 0 \ \dfrac{ft}{s}" is the final velocity of the projectile at a maximum height, "v_{y,i} = v_isin \theta" is the vertical component of the velocity of the projectile, "v_i = 1500 \ \dfrac{ft}{s}" is the initial velocity of the projectile, "\theta = 30^{\circ}" is the launch angle, "a = g =32 \ \dfrac{ft}{s^2}" is the acceleration due to gravity, "y_{max}" is the maximum height that the projectile can reach.

Then, from this equation we can find "y_{max}":


"y_{max} = \dfrac{(v_isin\theta)^2}{2g} = \dfrac{(1500 \ \dfrac{ft}{s} \cdot sin30^{\circ})^2}{2 \cdot 32 \ \dfrac{ft}{s^2}} = 8789 \ ft."

We can find the range of the projectile from the formula:


"R = \dfrac{v_i^2sin2\theta}{g} = \dfrac{(1500 \ \dfrac{ft}{s})^2 \cdot sin2\cdot 30^{\circ}}{32 \ \dfrac{ft}{s^2}} = 60892 \ ft."

Answer:

"y_{max} = 8789 \ ft, R = 60892 \ ft."